# How to Beat Nakamoto in the Race

### v1

@@ 2.5. Discussion
Traditional approaches often employ a view-based model incorporating elements of the universally composable security framework, executed in discrete time steps (pass2017analysis, ; gazi2020tight, ; dembo2020everything, ; canetti2020universally, ). However, modeling the views of individual nodes necessitates accounting for intricate interactions and dependencies, thereby incurring substantial mathematical and computational overhead. The tree-based mathematical model in Sec. 2.3 is considerably simpler yet self-contained. Rather than tracking each honest node’s view over time, we focus on the “weakest possible” view across all honest nodes. This weakest view, defined precisely in Definitions 3 and 4, includes only public chains. This tree-based model is sufficient to capture and evaluate all adversarial strategies that could lead to safety violations. In a nutshell, our insight on the weakest possible view eliminates the need to define individual views and their complex interactions. We maintain that our mathematical model is succinct, sufficient, and sound.
@@
We remark that all existing latency-security analyses of the Nakamoto consensus allow a wide range of non-causal attacks by giving the adversary foresight. For example, previous works (garay2015bitcoin, ; pass2017rethinking, ; dembo2020everything, ; cao2025security, ; guo2022bitcoin, ) all focus on the event where a greater number of certain types of blocks are mined compared to others during the confirmation period of interest. This is tantamount to allowing the adversary to optimally assemble blocks into chains with foresight of the timings of all future block arrivals. Those analyses generally do not achieve the ultimate fault tolerance for finite confirmation depths. An exception is (cao2025security, ), where the gap is closed by introducing a crucial causality requirement: Once the adversary falls behind the public height, honest mining power is concentrated on the public front, and the adversary can only catch up using freshly mined A-blocks, where no other blocks can be used to catch up with the public block.
@@ 2.6. Adversarial Policy (Attack)
In all but Sec. 7 and Appendix E of this paper, we restrict our attention to the safety of height 1 for clarity and simplicity. A well-known attack is the following:
@@
Under this private mining attack, all A-blocks form a chain, referred to as the A-chain. Height 1’s safety is violated as soon as the A-chain is both credible and consists of at least $k$ blocks. At this point, the A-chain confirms the first A-block to a depth of $k$ or more, while an H-block on height 1 is either already $k$-confirmed or will be once an H-block reaches height $k$.
@@
The first H-block to arrive at its height is called a jumper. The index of the highest jumper that arrives by $t$ is equal to $\min\{b\in\bm{H}_{t}\,|\,h_{b}\geq h_{d},\,\forall d\in\bm{H}_{t}\}$.
@@
For example, the blocktree in Fig. 1 contains two branches, branch 1 and branch 2. Branch 1 includes only block 1 at $t_{1}$. By $t_{6}$, branch 1 includes also blocks 3, 5, and 6 and is a sub-tree of the blocktree.
@@
Recall that by assumption the adversary can break ties arbitrarily for longest-chain fork choice. Thus it can make a credible chain the fork choice of all honest nodes. Formally we introduce the following notion:
@@
A snapshot of a successful bait-and-switch attack is given in Fig. 2 for PoW, assuming confirmation depth $k=4$. At $t_{8}$, chain 8 is on an unique highest branch, and it $k$-confirms block 1. Chain 7 is public. Chain 6 is a second highest branch and equal to public height 3, thus it is made the public fork choice at $t_{8}$. H-block 9 arrives by $t_{8}+\Delta$ and is used to extend chain 6, confirming block 2 to cause a safety violation.
@@
Figure 2. Example of a bait-and-switch attack under PoW consensus.
@@ 3. Main Results on Safety of Height 1
We recall that $a$ denotes the adversarial arrival rate, $h$ denotes the honest arrival rate, and $\Delta$ denotes the propagation delay upper bound. The total arrival rate is denoted by $\lambda=a+h$ and the fraction of adversarial arrivals is $\beta=a/\lambda$.
@@
We next evaluate the probability of violation achieved by the bait-and-switch attack.
@@
so the corresponding probability mass function (pmf) is obtained as
@@
(3) $\begin{split}&P_{W|L}(w|l)\\ &=\begin{cases}f_{1}(w,a\Delta),&\text{if }l<0\text{ or }w<l,\\ e^{-\lambda\Delta};&\text{if }l=0,w=l;\\ \int_{0}^{\Delta}f_{2}(t;l,a)e^{-\lambda(\Delta-t)}dt,&\text{if }l>0,w=l;\\ \int_{0}^{\Delta}\lambda e^{-\lambda s}f_{1}(w-l-1;a(\Delta-s))ds,&\text{if }l=0,w>l;\\ \int_{0}^{\Delta}f_{2}(t;l,a)\int_{0}^{\Delta-t}\lambda e^{-\lambda s}\\ \qquad\cdot f_{1}(w-l-1;a(\Delta-t-s))dsdt,&\text{if }l>0,w>l.\end{cases}\end{split}$
@@
then the probability of safety violation on height 1 is equal to
(7) $1-\sum_{i=0}^{k-1}\,\sum_{j=0}^{k-1-i}e(i)\left[P^{(1)}\times P^{(2)}\times\dots\times P^{(k)}\right]_{1,j+1},$
where $e(\cdot)$ is given by (2) in Appendix D and $P^{(1)},\dots,P^{(k)}$ denote some $(k+1)\times(k+1)$ transition probability matrices defined in (4) and (5), and $\cdot]_{1,j}$ takes the entry of a matrix on the first row and the $j$-th column.
@@
Figs. 3 and 4 plot the security-latency trade-off for Bitcoin ($\lambda=1/600$ blocks/second, $\Delta=10$ seconds) and Ethereum Classic (PoW, $\lambda=1/13$ blocks/second, $\Delta=2$ seconds), respectively. Computing the exact probabilities entails multiplying $k$ matrices of size $(k+1)\times(k+1)$, yielding a time complexity of $\mathcal{O}(k^{4})$. In Fig. 3, we plot the trade-off for Bitcoin given by (7) under three different adversarial fractions, with $\beta=0.4$, $0.25$, and $0.1$, respectively. Similarly, in Fig. 4, we plot the trade-off of Ethereum Classic (PoW) given by (7) under two different adversarial fractions. We also include upper and lower bounds based on results in (guo2022bitcoin, ) and (cao2025security, ) under the same parameters for reference.
@@ 4.1. State and Transition
Within the constraints of mining power and propagation delay, the adversary is allowed to continuously reallocate mining effort and arbitrarily delay block propagation. However, once we grant the enhanced adversary the privilege of placing each block at the moment of its arrival (cf. Sec. 2.2), the action simplifies: it suffices for the adversary to act only at discrete times—specifically, upon block arrivals. This naturally formulates the adversary’s task as a discrete-time control problem: Given the full history of the system up to the current arrival, where should the adversary place the new block to optimize its objective? Each adversarial action taken upon a block’s arrival transitions the system to a new state, which will be further updated at the next arrival. To formalize this setting, we introduce an MDP that captures the adversary’s decision-making problem.
The central challenge lies in crafting a state representation that is expressive enough to capture all relevant information, yet compact enough to render the MDP tractable. The most complete description of system state would include the full local views of all nodes. However, as discussed in Sec. 2.4, when analyzing safety violations, it suffices to consider only the public view—the weakest view shared by all honest nodes. Accordingly, we define the system state to consist of the following components: 1) the full blocktree, including the A/H classification of each block, 2) the tree’s public prefix, and 3) the arrival times of all non-public blocks.
@@
Upon a new arrival, the state is $(\mathcal{T},I)$, where $I\in\{A,H\}$. The adversary’s action consists of choosing a parent for the new block arrival. This action transitions the MDP to a new state $(\mathcal{T}^{\prime},\emptyset)$, where the blocktree $\mathcal{T}^{\prime}$ incorporates the newly added blocks and updates certain timers (per Def 16 in the case of an H-arrival). The $\emptyset$ indicates that this state immediately follows an action and precedes the next arrival. Upon the subsequent arrival, the system transitions to a new state $(\mathcal{T}^{\prime\prime},I^{\prime\prime})$, and so on.
@@ 4.2. MDP, Reward, and Markov Policy
This subsection defines state transition probabilities and a reward function for safety violations, and outlines the mathematical principles showing that a Markov policy satisfying the Bellman equation maximizes the probability of such violations. It also develops several lemmas which serve as the foundation for subsequent analysis.
@@
A decision rule $d_{i}$ prescribes how actions are chosen at decision epoch $i$. In general, a randomized, history-dependent rule defines a distribution over actions $Y_{i}$ based on the full history $H_{i}=(X_{1},Y_{1},\dots,X_{i-1},Y_{i-1},X_{i})$, where $X_{i}$ denotes the state at epoch $i$. A policy $\pi=(d_{1},d_{2},\dots)$ is a sequence of decision rules that induces a reward process $(X_{i},r(X_{i},Y_{i}))$, for $i=1,2,\dots$. A Markov definite decision rule is a special case where the action is deterministic and depends only on the current state, i.e., $Y=d(X)$. Such policies are particularly important, as they can be shown to achieve optimality in our problem.
@@
A safety violation occurs when a block is added to the blocktree, not upon a new arrival that is yet to be placed. With this in mind, we assign a unit reward the first time the system transitions into a state in $\mathscr{A}$:
@@
According to (puterman1994markov, , Theorem 7.1.3), the function $v^{*}$ must satisfy the Bellman optimality equation:
@@ A Markov definite decision rule $d$ is conserving if for all $s\in\mathcal{S}$,
Within the ultimate fault tolerance (6), the probability of violating height-1 safety under the $k$-confirmation commitment rule vanishes as $k\to\infty$. Since H-block heights necessarily increase without bound over time, it follows that the probability of a first safety violation occurring after $n$ arrivals also vanishes as $n\to\infty$ under condition (6). Consequently, a stationary policy $\pi=(d,d,\dots)$ formed by repeating a conserving decision rule $d$ is an equalizing policy in the sense of (puterman1994markov, , Eq. 7.0.950), i.e.,
@@
With the preceding technical conditions verified, we invoke (puterman1994markov, , Theorem 7.1.7 b) to conclude the following:
@@ 4.3. Lemmas for Comparing States
The remaining challenge is to determine an optimal action to take upon each arrival. A Markov definite decision rule uniquely determines the next state, and an optimal action is one that transitions the system to a state with the highest value. For $s\notin\mathscr{A}$, the right-hand side of (13) is equivalent to $\max_{\alpha\in\mathcal{A}_{s}}\{V(u)\}$, where $u$ denotes the resulting next state from action $\alpha$. Therefore, finding an optimal decision rule $d$ boils down to comparing the values of all feasible next states resulting from admissible actions.
@@ 4.4. Simplification of State Representation
If states $s$ and $s^{\prime}$ have identical blocktrees except that, for every block, the timer in $s$ is greater than or equal to the timer in $s^{\prime}$, then $s\succeq s^{\prime}$.
The proof is given in Appendix A.5.
To simplify the state $s=(\mathcal{T},I)$, we partition the entire blocktree $\mathcal{T}$ into distinct branches according to Definition 13. The number of branches is equal to the number of height-1 blocks in the blocktree. We demonstrate in the following lemma, which is proved in Appendix A.6, that only the two highest of these branches need to be recorded in the state representation and that it suffices to store only the minimum block timer on each height.
@@ 4.5. Public Height, Violation, and Transitions
where we use the fact that the timer on height $(m\land d)-1$ is set to zero. By definition of $\mathscr{A}$ in Sec. 4.2, we have $s\in\mathscr{A}$ if and only if
@@
In any given state $m,d,n,(l_{m\land d},\dots,l_{d}),I]$, we explicitly describe the state transition rule in Table 2 for all allowable actions, depending on the conditions of the state under which the actions are applied. For $I=\emptyset$, there is no block to place and the “action” is to await an arrival, so we let $\emptyset$ denote the action and $A_{s}=\{\emptyset\}$ in this case. The subsequent arrival transitions the state probabilistically as described in Table 2. For $I\in\{A,H\}$, the action deterministically transitions the system to the next state. For $I=H$, the H-block can be placed on either branch strictly higher than the public height, with the action set $A_{s}=\{\overline{P+1},\dots,\overline{n+1}\}\cup\{\underline{P+1},\dots,\underline{m+1}\}.$ For $I=A$, the A-block can be placed on any height, so the action set $A_{s}=\{\overline{1},\dots,\overline{n+1}\}\cup\{\underline{1},\dots,\underline{m+1}\}.$
@@
An example of the trajectory of state transitions is as follows: Starting from state $1,2,3,(\Delta)]$, an A-arrival occurs after time $\frac{1}{3}\Delta$, transitioning the state to $1,2,3,(\frac{2}{3}\Delta),A]$. With $\underline{2}$ taken, the system transitions to state $2,2,3,(\frac{2}{3}\Delta)]$. An H-arrival occurs $10\Delta$ later, transitioning the state to $[2,2,3,(0),H]$. The adversary elects to take action $\underline{3}$, causing a transition to $3,3,3,(\Delta)]$. We have $3\geq\max\{3,2\}$, thus state $3,3,3,(\Delta)]\in\mathscr{A}$ is a violation if the depth $k\leq 3$.

### v2

@@ 2.5. Discussion
Traditional approaches often employ a view-based system incorporating elements of the universally composable security framework, executed in discrete time steps (Pass et al., 2017; Gaži et al., 2020; Dembo et al., 2020; Canetti, 2020). However, modeling the views of individual nodes necessitates accounting for intricate interactions and dependencies, thereby incurring substantial mathematical and computational overhead. The tree-based mathematical model in Sec. 2.3 is considerably simpler yet self-contained. Rather than tracking each honest node’s view over time, we focus on the “weakest possible” view across all honest nodes. This weakest view, defined precisely in Definitions 3 and 4, includes only public chains. This tree-based model is sufficient to capture and evaluate all adversarial strategies that could lead to safety violations. In a nutshell, our insight on the weakest possible view eliminates the need to define individual views and their complex interactions. We maintain that our mathematical model is succinct, sufficient, and sound.
@@
We remark that all existing latency-security analyses of the Nakamoto consensus allow a wide range of non-causal attacks by giving the adversary foresight. For example, previous works (Garay et al., 2015; Pass and Shi, 2017a; Dembo et al., 2020; Cao and Guo, 2025; Guo and Ren, 2022) all focus on the event where a greater number of certain types of blocks are mined compared to others during the confirmation period of interest. This is tantamount to allowing the adversary to optimally assemble blocks into chains with foresight of the timings of all future block arrivals. Those analyses generally do not achieve the ultimate fault tolerance for finite confirmation depths. An exception is (Cao and Guo, 2025), where the gap is closed by introducing a crucial causality requirement: Once the adversary falls behind the public height, honest mining power is concentrated on the public front, and the adversary can only catch up using freshly mined A-blocks, where no other blocks can be used to catch up with the public block.
Our mathematical model does not incorporate cryptographic mechanisms. For instance, arrival processes are exogenous rather than resulting from cryptographic hash computations. Consequently, we do not introduce a cryptographic security parameter, such as the probability of hash collisions, as used in related works to quantify block safety (see e.g., (Pass et al., 2017; Garay et al., 2020)). This omission does not compromise practicality, as the probability of a block safety violation of interest is orders of magnitude higher than that of a cryptographic failure.
@@ 2.6. Adversarial Policy (Attack)
In all but Sec. 7, we restrict our attention to the safety of height 1 for clarity and simplicity. A well-known attack is the following:
@@
Under this private mining attack, all A-blocks form a chain, referred to as the A-chain. Height 1’s safety is violated as soon as the A-chain is both credible and consists of at least $k$ blocks. At this point, the A-chain confirms the first A-block at a depth of $k$ or more, while an H-block at height 1 is either already $k$-confirmed or will be once an H-block reaches height $k$.
@@
The first H-block to arrive at its height is called a jumper.
@@
For example, the blocktree in Fig. 1 contains two branches, branches 1 and 2. Branch 1 includes only block 1 at $t_{1}$. By $t_{6}$, branch 1 includes also blocks 3, 5, and 6.
@@
Recall that the adversary can break ties arbitrarily for longest-chain fork choice. Thus it can make a credible chain the fork choice of all honest nodes. Formally we introduce the following notion:
@@
A snapshot of a successful bait-and-switch attack is illustrated in Fig. 2, assuming confirmation depth $k=4$. At $t_{8}$, chain 8 is on a unique highest branch, and it $k$-confirms block 1. Chain 7 is public. Chain 6 is a second highest branch and equal to public height 3, thus it is made the public fork choice at $t_{8}$. H-block 9 arrives by $t_{8}+\Delta$ and is used to extend chain 6, confirming block 2 to cause a safety violation.
@@
Figure 2. Example of a successful bait-and-switch attack.
@@ 3. Main Results on Safety of Height 1
We recall that $a$ represents the adversarial arrival rate, $h$ represents the honest arrival rate, and $\Delta$ represents the upper bound on network delays. In the remainder of this paper, we denote the total arrival rate as $\lambda=a+h$, and the fraction of adversarial arrivals as $\beta=a/\lambda$.
@@
We shall develop a proof of Theorem 1 in Sec. 6. We next evaluate the probability of violation achieved by the bait-and-switch attack.
@@
so the corresponding probability mass function (pmf) is
@@
(3) $\begin{split}&P_{W|L}(w|l)\\ &=\begin{cases}f_{1}(w,a\Delta),&\text{if }l<0\text{ or }w<l;\\ e^{-\lambda\Delta},&\text{if }l=0,w=l;\\ \int_{0}^{\Delta}f_{2}(t;l,a)e^{-\lambda(\Delta-t)}dt,&\text{if }l>0,w=l;\\ \int_{0}^{\Delta}\lambda e^{-\lambda s}f_{1}(w-l-1;a(\Delta-s))ds,&\text{if }l=0,w>l;\\ \int_{0}^{\Delta}f_{2}(t;l,a)\int_{0}^{\Delta-t}\lambda e^{-\lambda s}\\ \qquad\cdot f_{1}(w-l-1;a(\Delta-t-s))dsdt,&\text{if }l>0,w>l.\end{cases}\end{split}$
@@
then the probability of safety violation at height 1 is equal to
(7) $1-\sum_{i=0}^{k-1}\,\sum_{j=1}^{k-i}e(i)\left[P^{(1)}\times P^{(2)}\times\dots\times P^{(k)}\right]_{1,j},$
where $e(\cdot)$ is given by (2) and $P^{(1)},\dots,P^{(k)}$ denote some $(k+1)\times(k+1)$ transition probability matrices defined in (4) and (5), and $\cdot]_{1,j}$ takes the entry of a matrix on the first row and the $j$-th column.
Theorem 2 is proved in Appendix D.
@@
Figs. 3 and 4 plot the security-latency trade-off for Bitcoin ($\lambda=1/600$ blocks/second, $\Delta=10$ seconds) and Ethereum Classic (PoW, $\lambda=1/13$ blocks/second, $\Delta=2$ seconds), respectively. Computing the exact probabilities entails multiplying $k$ matrices of size $(k+1)\times(k+1)$, yielding a time complexity of $\mathcal{O}(k^{4})$. In Fig. 3, we plot the trade-off for Bitcoin given by (7) under three different adversarial fractions, with $\beta=0.4$, $0.25$, and $0.1$, respectively. Similarly, in Fig. 4, we plot the trade-off of Ethereum Classic (PoW) given by (7) under three different adversarial fractions. We also include upper and lower bounds based on results in (Guo and Ren, 2022) and (Cao and Guo, 2025) under the same parameters for reference.
@@ 4.1. State and Transition
Within the constraints of mining power and propagation delay, the adversary is allowed to continuously reallocate mining effort and arbitrarily delay block propagation. However, once we grant the enhanced adversary the privilege of placing each block at the moment of its arrival (cf. Sec. 2.2), the action simplifies: it suffices for the adversary to act only at discrete times—specifically, upon block arrivals. This naturally formulates the adversary’s task as a discrete-time control problem: Given the full history of the system up to the current arrival, where should the adversary place each new block to optimize its objective? Each adversarial action taken upon a block’s arrival transitions the system to a new state, which will be further updated at the next arrival. To formalize this setting, we introduce an MDP that captures the adversary’s decision-making problem.
The central challenge lies in crafting a state representation that is expressive enough to capture all relevant information, yet compact enough to render the MDP tractable. The most complete description of system state would include the full local views of all nodes. However, as discussed in Sec. 2.4, when analyzing safety violations, it suffices to consider only the public view—the weakest view shared by all honest nodes. Accordingly, we first consider defining the system state to consist of the following components: 1) the full blocktree, including the A/H classification of each block, 2) the tree’s public prefix, and 3) the arrival times of all non-public blocks.
@@
Upon a new arrival, the state is $(\mathcal{T},I)$, where $I\in\{A,H\}$. The adversary’s action consists of choosing a parent for the new block arrival. This action transitions the MDP to a new state $(\mathcal{T}^{\prime},\emptyset)$, where the blocktree $\mathcal{T}^{\prime}$ incorporates the newly added blocks and updates certain timers (per Definition 16 in the case of an H-arrival). The $\emptyset$ in $(\mathcal{T}^{\prime},\emptyset)$ indicates that this state immediately follows an action and awaits the next arrival. Upon the subsequent arrival, the system transitions to a new state $(\mathcal{T}^{\prime\prime},I^{\prime\prime})$, and so on.
@@ 4.2. MDP, Reward, and Markov Policy
This subsection defines state transition probabilities and a reward function for safety violations, and outlines the mathematical principles showing that a Markov policy maximizes the probability of such violations. It also develops several lemmas which serve as the foundation for subsequent analysis.
@@
A decision rule $d_{i}$ prescribes how actions are chosen at decision epoch $i$. In general, a randomized, history-dependent rule defines a distribution over actions $Y_{i}$ based on the full history $H_{i}=(X_{1},Y_{1},\dots,X_{i-1},Y_{i-1},X_{i})$, where $X_{i}$ denotes the state at epoch $i$. A policy $\pi=(d_{1},d_{2},\dots)$ is a sequence of decision rules that induces a reward process $(X_{i},r(X_{i},Y_{i}))$, for $i=1,2,\dots$. A Markov definite decision rule is a special case where the action is deterministic and depends only on the current state, i.e., $Y_{i}=d(X_{i})$. Such policies are particularly important, as they can be shown to achieve optimality in our problem.
@@
A safety violation occurs only when some blocks is placed onto the blocktree, not upon a new arrival that is yet to be placed. With this in mind, we assign a unit reward the first time the system transitions into a state in $\mathscr{A}$:
@@
According to (Puterman, 1994, Theorem 7.1.3), the function $v^{*}$ must satisfy the Bellman optimality equation:
@@ A Markov definite decision rule $d$ is conserving if for all $s\in\mathcal{S}$,
Within the ultimate fault tolerance (6), the probability of violating height-1 safety under the $k$-confirmation commitment rule vanishes as $k\to\infty$. Since H-block heights necessarily increase without bound over time, it follows that the probability of a first safety violation occurring after $n$ arrivals also vanishes as $n\to\infty$ under condition (6). Consequently, a stationary policy $\pi=(d,d,\dots)$ formed by repeating a conserving decision rule $d$ is an equalizing policy in the sense of (Puterman, 1994, Eq. 7.0.950), i.e.,
@@
With the preceding technical conditions verified, we invoke (Puterman, 1994, Theorem 7.1.7 b) to conclude the following:
@@ 4.3. Lemmas for Comparing States
The remaining challenge is to determine an optimal action to take upon each arrival. A Markov definite decision rule uniquely determines the next state, and an optimal action is one that transitions the system to a state with the highest value. For $s\notin\mathscr{A}$, the right-hand side of (13) is equivalent to $\max_{\alpha\in\mathcal{A}_{s}}\{{r(s,\alpha)+}V(u)\}$, where $u$ denotes the resulting next state from action $\alpha$. Therefore, if no action transitions the state $s$ to violation in one step (i.e., $r(s,\alpha)=0$ for all $\alpha\in\mathcal{A}_{s}$), finding an optimal action boils down to comparing the values of all feasible next states.
@@ 4.4. Simplification of State Representation
If states $s$ and $s^{\prime}$ have identical blocktrees except that, for every block, its timer in $s$ is greater than or equal to its timer in $s^{\prime}$, then $s\succeq s^{\prime}$.
[…]
Figure 5. An example for Lemma 5, where $s\succeq s^{\prime}$.
Lemma 5 is proved in Appendix A.5. An example illustrating the lemma is provided in Fig. 5.
To simplify the state $s=(\mathcal{T},I)$, we partition the entire blocktree $\mathcal{T}$ into distinct branches according to Definition 13. The number of branches is equal to the number of height-1 blocks in the blocktree. It can be shown that only the two highest of these branches need to be recorded in the state representation and that it suffices to store only the minimum block timer on each height. We demonstrate this in the following lemma, which is proved in Appendix A.6.
@@ 4.5. Public Height, Violation, and Transitions
where we use the fact that the timer at height $(m\land d)-1$ is set to zero. By definition of $\mathscr{A}$ in Sec. 4.2, the state is in violation, i.e., $s\in\mathscr{A}$, if and only if
@@
In any given state $m,d,n,(l_{m\land d},\dots,l_{d}),I]$, we explicitly describe the state transition rule in Table 2 for all allowable actions, depending on the conditions of the state under which the actions are applied. For $I=\emptyset$, there is no block to place and the “action” is to await an arrival, so we let $\emptyset$ denote the action and $A_{s}=\{\emptyset\}$ in this case. The subsequent arrival transitions the state probabilistically as described in Table 2. For $I\in\{A,H\}$, the action deterministically transitions the system to the next state. For $I=H$, the H-block can be placed on either branch strictly higher than the public height, with the action set $A_{s}=\{\overline{P+1},\dots,\overline{n+1}\}\cup\{\underline{P+1},\dots,\underline{m+1}\}.$ For $I=A$, the A-block can be placed at any height, so the action set $A_{s}=\{\overline{1},\dots,\overline{n+1}\}\cup\{\underline{1},\dots,\underline{m+1}\}.$
@@
We illustrate the evolution of the blocktree with associated timers using an example in Fig. 6. The corresponding arrivals, actions, and state transitions are as follows: Starting from state $1,2,3,(0,\Delta)]$, an A-arrival occurs after time $\frac{1}{3}\Delta$, transitioning the state to $1,2,3,(0,\frac{2}{3}\Delta),A]$. With $\underline{2}$ taken, the system transitions to state $2,2,3,(\frac{2}{3}\Delta)]$. An H-arrival occurs ${2}\Delta$ later, transitioning the state to $[2,2,3,(0),H]$. The adversary elects to take action $\underline{3}$, causing a transition to $3,3,3,(\Delta)]$. We have $3\geq\max\{3,2\}$, thus state $3,3,3,(\Delta)]\in\mathscr{A}$ is a violation if the depth $k\leq 3$.
[…]
Figure 6. Example of state transition.